OpenStudy (sanaelshazly):

The potential difference across two capacitors in series is 210. V. the capacitances are C1=3400. mf. and C2=3940 mf. a) what is the capacitance of this pair of capacitors ? Ceq=............ mf.

OpenStudy (sanaelshazly):

it tells me that the answer is wrong !

OpenStudy (anonymous):

sorry i misread it .....working on it

OpenStudy (sanaelshazly):

:D ThnQQ !

OpenStudy (anonymous):

is it 1825.06 mf?

OpenStudy (sanaelshazly):

it tells me that it's wrong answer ! again ! =/

OpenStudy (anonymous):

have you provided the correct details? 1/c1 +1/c2

OpenStudy (anonymous):

thats how we calculate capacitance for 2 capacitors in series.

OpenStudy (sanaelshazly):

(c1*c2)/(c1+c2)

OpenStudy (anonymous):

thats correct too product/sum 1/c1 +1/c2 = c1*c2/c1+c2 ;)

OpenStudy (sanaelshazly):

i didn't get it actually !

OpenStudy (anonymous):

whats the correct answer?

OpenStudy (sanaelshazly):

it didnt tell me !!

OpenStudy (sanaelshazly):

it is telling me try again !

OpenStudy (anonymous):

try this 7340

OpenStudy (sanaelshazly):

7340 it's right :D b) what is the charge on each capacitor ? q1=..........C q2=..........C c) what is the potential difference across each capacitor ? V1=..........V V2=.........V d) what is the total energy stored by the capacitors ? U=........J

OpenStudy (anonymous):

okay so i did this. firstly i just calculated the charge on each capacitor for C1 Q=CV Q=.714 for C2 Q=CV Q=.8274 Total Q = Q1 + Q2 = 1.5414 Now we have total Q put it in equation Q=CV 1.5414=C*210 C=7340mF

OpenStudy (sanaelshazly):

ThnQQ !!

OpenStudy (anonymous):

Energy = 1/2 *C*V^2 =1/2 * 7340mf*210^2 =161.847J